OpenStudy (anonymous):

2(6z-1)-3(z+6)=4(z+1) The solution z =?

OpenStudy (anonymous):

12z-2-3z-18=4z+4 5z=24 z=24/5

OpenStudy (anonymous):

thank you I totally understand this equation and how to process it to get that answer.

OpenStudy (anonymous):

the answer is z=24/5